You are considering the purchase of a new stock. The stock is forecasted to pay a dividend next year (D1) of $3.98. In addition, you forecast that the firm will have a stable growth rate of 4.5% for the foreseeable future. The current risk-free rate of return is 4.4%. The expected return on the market is 7.8% and the standard deviation for the market is 18%. The stock has a correlation to the market of 0.29. Finally, the stock has a standard deviation of 44%.

Given this information, what is the value of this stock? (Hint: Use the constant growth pricing model)
Answer #1

Beta of Stock =Correlation * Standard Deviation of Stock/Standard Deviation of Market =0.29*44%/18% =0.7089
CAPM using cost of equity =Risk free rate+beta*(Expected Return -Risk free rate) =4.4%+0.7089*(18%-4.4%) =14.04089%
Value of the Stock price =Dividend 1/(Required rate -growth) =3.98/(14.04089%-4.5%) =41.72